OpenStudy (anonymous):

The half life for the decay of radium is 1620 years. What is the rate constant for this first order process?

OpenStudy (aaronq):

first order: \[t _{1/2}=\frac{ \ln2 }{ k }\] t1/2= half life k = rate constant